Calculate the volume occupied at 0 degree Celsius by 100g of liquid water & by 100g of Ice?&
Calculate the percentage increased in volume when 100g of water freezes at 0 degree Celsius?&
Calculate the concentration of water (in M) at 0 degree Celsius?
1
Expert's answer
2011-10-06T09:52:15-0400
Calculate the volume occupied at 0 degree Celsius by 100g of liquid water
& by 100g of Ice?
Calculate the percentage increased in volume when 100g of water freezes at
0 degree Celsius?
Calculate the concentration of water (in M) at 0 degree Celsius?

The density of liquid water at 0 degree Celsius is 1 g/ml. So, the volume of 100g liquid water at 0 degree Celsius is 100ml.
The density of ice is 0,9 g/ml. So, the volume of 100g of ice is (100g)/(0.9g/ml) = 111ml.
The percentage increased in volume when 100g of water freezes at 0 degree Celsius is (111ml/100ml)*100% = 111%.
Let`s calculate the concentration of water at 0 degree Celsius.
M(H2O) = 2M(H) +M(O) = 2·1+16 = 18 (g/mol)
(100g)/(18g/mol) = 5.55 mol
So, the concentration of water at 0 degree Celsius is (5.55mol)/(100ml) = 0.0555 mol/ml.